# SIC: Similarity-Based Interpretable Image Classification with Neural Networks

### v2

@@ 1 Introduction
As an alternative, inherently interpretable models provide explanations by design. For instance, logistic regression can be considered an inherently interpretable model since evaluating its coefficients directly explains the effect of each input feature towards the prediction. Translating these properties to complex functions implemented by neural networks was further fueled by the argument that if a technique is able to summarize the decision-making of a black box without any error, the black box could be replaced by an inherently interpretable model in the first place [26]. As a result, a growing research community is dedicated to developing inherently interpretable models, e.g., for decision-critical tasks like medicine [4, 16, 32].
@@
  • We empirically evaluate SIC and its explanations on three tasks (fine-grained single-label image classification on Stanford Dogs, pathology prediction on RSNA, multi-label image classification on Pascal VOC), and three architectures with different numbers of parameters (DenseNet121: 8M, ResNet50: 26M, a hybrid vision transformer: 81M).
@@ 2 Background
$p(y=c\mid\mathcal{I})=\sum_{k}\frac{\exp\left(-\left\|\mathcal{F}_{\theta}(\mathcal{I})-v_{k}^{c}\right\|_{2}^{2}\right)}{\sum_{k^{\prime}=1}^{N_{s}}\sum_{c^{\prime}=1}^{C}\exp\left(-\left\|\mathcal{F}_{\theta}(\mathcal{I})-v_{k^{\prime}}^{c^{\prime}}\right\|_{2}^{2}\right)}.$
@@ 3 Methods
We propose SIC, which combines the strengths of B-cos networks and the Nadaraya-Watson head, to obtain an inherently interpretable neural network for similarity-based classification that provides faithful local and global explanations with a focus on minimizing the number of explanations requiring review. SIC achieves this by extracting class-representative support vectors from training images for similarity-based classification and by leveraging the B-cos transform to faithfully explain its decision.
@@ 3.1 Architecture
The evidence predictor $\mathcal{E}$ consists of a function $\oplus:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}_{\geq 0}$ that transforms the real-valued vectors $f$ into non-negative vectors $f^{+}=\oplus(f)$, and a similarity measure that computes similarities between support vectors $v_{i}^{c}\in F^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}$ and $f^{+}$ as $sim(f^{+},v_{i}^{c})\in\mathbb{R}$, with $\Omega(F^{c})=N_{s}$ and $i$ denoting the $i$-th support vector of class $c$. Logits are computed for class $c$ as the sum of its similarity scores $\mu_{c}=b+\sum_{v_{i}^{c}}\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}}$, where $b$ represents a fixed bias term, $i=1,\dots,N_{s}$ indexes the support samples, and $\mathcal{T}$ is a temperature to improve convergence by scaling the logits’ magnitude appropriately. Examples of suitable transformation functions $\oplus$ include the exponential function, absolute value, and ReLU.
@@ 3.2 Training and Optimization
Following Wang and Sabuncu [31], we randomly sample the training set to extract $N_{s}$ support vectors $v_{k}^{c}$ per class to encourage the model to form distinct, class-specific clusters in latent space. After every epoch, we compute sets of latent vectors $\hat{V}^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}_{k=1}^{N}$, which we each cluster using k-means to extract centroids $\gamma_{i}^{c}$. Then, we replace each centroid $\gamma_{i}^{c}$ with the closest latent vector to yield support vectors from the same class, defined as $v_{i}^{c}=\arg\min_{f^{+}_{k}\mid y_{k}=c}\|f^{+}_{k}-\gamma_{i}^{c}\|_{2}$, which facilitates case-based reasoning similar to k-nearest-neighbors, as only features descriptive of the class can increase the probability of that class by definition of SIC.
@@ 4 Experiments and Results
First, we compare the performance of SIC in terms of the (balanced) accuracy to black-box neural networks and established inherently interpretable neural networks, namely ProtoPNet [8], a BagNet17 [6], the NW-Head [31], and B-Cos neural networks [7] in Fig. 2 and Tab. A.2. Compared to black-box versions of the backbones, SIC performs on par on Pascal VOC and RSNA ($\geq+0.29\%$ and $-0.65\%$ Acc, respectively), and worse on Stanford Dogs ($-3.7\%$ Acc). Overall, ProtoPNet and BagNets perform worse than all competing methods, and SIC performs better than nine out of 14 models.
@@ 5 Discussion and Conclusion
We demonstrated that its intuitive global and local explanations not only facilitate effective model debugging for developers but also simplify the understanding of its decision-making process for end-users.
@@ A.3 Implementation and Training Details
All backbones are extended by a linear projection layer to extract feature vectors $\in\mathbb{R}^{d}=128$. The $\oplus$ function is implemented as a ReLU activation in the evidence predictor $\mathcal{E}$. In multi-label classification settings, only training samples with a single class label are considered for support labels.
@@ A.4.4 RSNA - DenseNet121
Figure A.16: t-SNE plot of the latent vectors of the RSNA training set. Red denotes samples of the class ”opacity” and blue the class ”no opacity”. Support vector 0 and support vector 3 are very close in the projected latent visualization.

### v3

@@ 1 Introduction
As an alternative, inherently interpretable models provide explanations by design. For instance, logistic regression can be considered an inherently interpretable model since evaluating its coefficients directly explains the effect of each input feature towards the prediction. Translating these properties to complex functions implemented by neural networks was further fueled by the argument that if a technique is able to summarize the decision-making of a black-box without any error, the black-box could be replaced by an inherently interpretable model in the first place [26]. As a result, a growing research community is dedicated to developing inherently interpretable models, e.g., for decision-critical tasks like medicine [4, 16, 33].
@@
  • We empirically evaluate SIC and its explanations on three tasks (fine-grained single-label image classification on Stanford Dogs, pathology prediction on RSNA, multi-label image classification on Pascal VOC), and three architectures with different numbers of parameters (DenseNet121: $\sim$8M, ResNet50: $\sim$26M, a hybrid vision transformer: $\sim$81M).
@@ 2 Background
$p(y=c\mid\mathcal{I})=\displaystyle\sum_{k}\frac{\exp\left(-\left\|\mathcal{F}_{\theta}(\mathcal{I})-v_{k}^{c}\right\|_{2}^{2}\right)}{\displaystyle\sum_{k^{\prime}=1}^{N_{s}}\displaystyle\sum_{c^{\prime}=1}^{C}\exp\left(-\left\|\mathcal{F}_{\theta}(\mathcal{I})-v_{k^{\prime}}^{c^{\prime}}\right\|_{2}^{2}\right)}.$
@@ 3 Methods
We propose SIC, which combines the strengths of B-cos networks and the Nadaraya-Watson head, to obtain an inherently interpretable neural network for similarity-based classification that provides faithful local and global explanations with a focus on minimizing the number of explanations requiring review. SIC achieves this by extracting class-representative support vectors from training images for similarity-based classification and by leveraging the B-cos transform to faithfully explain its decision. Our implementation is available at www.github.com/ai-med/SIC.
@@ 3.1 Architecture
The evidence predictor $\mathcal{E}$ consists of a function $\oplus:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}_{\geq 0}$ that transforms the real-valued vectors $f$ into non-negative vectors $f^{+}=\oplus(f)$, and a similarity measure that computes similarities between support vectors $v_{i}^{c}\in F^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}$ and $f^{+}$ as $sim(f^{+},v_{i}^{c})\in\mathbb{R}$, with $\Omega(F^{c})=N_{s}$ and $i$ denoting the $i$-th support vector of class $c$. Logits are computed for class $c$ as the sum of its similarity scores $\mu_{c}=b+\sum_{v_{i}^{c}}\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}}$, where $b$ represents a fixed bias term, $i=1,\dots,N_{s}$ indexes the support samples, and $\mathcal{T}$ is a temperature to improve convergence by scaling the logits’ magnitude appropriately.
@@ 3.2 Training and Optimization
Following Wang and Sabuncu [31] (“closest cluster”), we randomly sample the training set to extract $N_{s}$ support vectors $v_{k}^{c}$ per class to encourage the model to form distinct, class-specific clusters in latent space. After every epoch, we compute sets of latent vectors $\hat{V}^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}_{k=1}^{N}$, which we each cluster using k-means to extract centroids $\gamma_{i}^{c}$. Then, we replace each centroid $\gamma_{i}^{c}$ with the closest latent vector to yield support vectors from the same class, defined as $v_{i}^{c}=\arg\min_{f^{+}_{k}\mid y_{k}=c}\|f^{+}_{k}-\gamma_{i}^{c}\|_{2}$, which facilitates case-based reasoning similar to k-nearest-neighbors, as only features descriptive of the class can increase the probability of that class by definition of SIC.
@@ 4 Experiments and Results
First, we compare the performance of SIC in terms of the accuracy (balanced accuracy on RSNA) to black-box neural networks and established inherently interpretable neural networks, namely ProtoPNet [8], a BagNet17 [6], the NW-Head [31], and B-Cos neural networks [7] in Fig. 2 and Tab. A.2. Compared to black-box versions of the backbones, SIC performs on par on Pascal VOC and RSNA ($\geq+0.29\%$ and $-0.65\%$ Acc, respectively), and worse on Stanford Dogs ($-3.7\%$ Acc). Overall, ProtoPNet and BagNets perform worse than all competing methods, and SIC performs better than nine out of 14 models.
@@ 5 Discussion and Conclusion
Our results suggest that its intuitive global and local explanations may not only facilitate effective model debugging for developers but may also simplify the understanding of its decision-making process for end-users.
@@ A.3 Implementation and Training Details
*: optimized with training recipe published in [9] instead of [8].
@@
All backbones are extended by a linear projection layer to extract feature vectors $\in\mathbb{R}^{d}=128$. The $\oplus$ function is implemented as a ReLU activation in the evidence predictor $\mathcal{E}$, and the similarity function “sim” is implemented by a B-cos linear layer. In multi-label classification settings, only training samples with a single class label are considered for support labels.
@@ A.3.1 FunnyBirds - ResNet50
A.3.2 CUB-200-2011 - ResNet50
We trained SIC on the official splits of CUB-200-2011 [32] (cropped ROI), achieving an accuracy of 79.8% (cmp. ProtoPNet 76.1-80.2% in Tab. 1 (top) [8]).
@@ A.4.4 RSNA - DenseNet121
Figure A.16: t-SNE plot of the latent vectors of the RSNA training set. Red denotes samples of the class "opacity" and blue the class "no opacity". Support vector 0 and support vector 3 are very close in the projected latent visualization.